# MDPs with a State Sensing Cost

### v1

@@ Abstract
In many practical sequential decision-making problems, tracking the state of the environment incurs a sensing/communication/computation cost. In these settings, the agent’s interaction with its environment includes the additional component of deciding when to sense the state, in a manner that balances the value associated with optimal (state-specific) actions and the cost of sensing. We formulate this as an expected discounted cost Markov Decision Process (MDP), wherein the agent incurs an additional cost for sensing its next state, but has the option to take actions while remaining ‘blind’ to the system state.
We pose this problem as a classical discounted cost MDP with an expanded (countably infinite) state space. While computing the optimal policy for this MDP is intractable in general, we bound the sub-optimality gap associated with optimal policies in a restricted class, where the number of consecutive non-sensing (a.k.a., blind) actions is capped. We also design a computationally efficient heuristic algorithm based on policy improvement, which in practice performs close to the optimal policy. Finally, we benchmark against the state of the art via a numerical case study.
Machine Learning, ICML
@@ 1 Introduction
Markov Decision Processes (MDPs) constitute an important framework for capturing the sequential interaction between an agent and an adaptive environment that ‘responds’ to the actions of the agent. The classical framework of an MDP is that at each time $t,$ the agent sees the state, say $X_{t},$ of the environment, and takes an action $A_{t}.$ This action, in turn, generates a feedback signal (reward/cost) that depends on the state-action pair $(X_{t},A_{t}),$ and triggers a random, action-dependent, Markovian transition in the state.
However, in many applications, ‘seeing’ the current state involves a cost. For example:
$\bullet$ In healthcare applications, there is often a monetary and/or delay cost associated with sensing the state of the patient during an ongoing intervention; e.g., white blood cell monitoring during anti-HIV drug administration, laboratory tests for ICU patients (Ernst et al., 2006; Jha et al., 2009).
$\bullet$ Applications on mobile phones must pay an energy cost to sense the location/motion/environment of the user, which must be balanced with the goal of enhancing user experience (Wang et al., 2010). Similarly, in a wireless sensor network, sensing the network state might require turning on battery-operated (and therefore energy-constrained) sensors.
$\bullet$ In remote surveillance applications, there is a communication cost to be incurred in transmitting state information (in the form of images/video) to the controller.
$\bullet$ In distributed sensing applications, there is often an additional computational cost in aggregating the sensor readings to obtain the state of the system.
$\bullet$ In robotics, where an autonomous robot is engaged in a certain task, sensing the surroundings (i.e., the state) of the robot might induce a latency (cost) in task completion.
This motivates the incorporation of a cost of state sensing, as well as a model for opportunistic state sensing, into the MDP framework. However, such an incorporation introduces technical challenges. In particular, if the agent chooses not to sense the system state (while continuing to interact with the environment), it is left with a belief distribution over the system state. Incorporating this belief into the MDP formulation induces a state space explosion, which makes the decision problem intractable.
@@
While the ‘augmented’ MDP of interest admits a stationary Markov policy, it is computationally intractable given the infinite state space. The goal of this paper is to design tractable, near-optimal algorithms for solving it. Our key contributions are as follows:
@@
$\bullet$ We derive computable bounds on the suboptimality associated with the optimal policies corresponding to the truncated MDPs.
@@
$\bullet$ We prove that if the state sensing cost is less than a certain computable threshold, then always sensing is optimal.
$\bullet$ We propose a policy-improvement-based heuristic algorithm, which selectively searches for improving blind action sequences, that is near-optimal in practice.
$\bullet$ Finally, we conduct an extensive numerical case study, comparing the proposed planning algorithms with the state of the art.
@@
It is important to note that the proposed formulation can be posed as a Partially Observable Markov Decision Processes (POMDP) (indeed, we do benchmark the proposed algorithms against state of the art POMDP solvers; see Section 5). However, in doing so, one loses the specific problem structure that arises in the opportunistic state sensing formulation, which we seek to exploit here for computational tractability; POMDPs are known to be intractable in general (Papadimitriou & Tsitsiklis, 1987).
The remainder of this paper is organized as follows. After surveying the literature below, we formulate our sensing-cost-incorporated MDP in Section 2. In Section 3, we propose a heuristic algorithm for this MDP. Next, in Section 4, we prove sufficient conditions for the optimality of a certain class of policies, and bound their suboptimality when they do not meet these conditions. Numerical case studies are presented in Section 5, and we conclude in Section 6. Proofs of our analytical results have been omitted from the main body of the paper due to space constraints; these can be found in the appendix.
Related Literature: In light of the preceding discussion on POMDPs, we focus here only on the (few) papers that study an explicit cost-sensing formulation identical (or similar) to ours.
@@
Formulations equivalent to the one in the present paper have been analysed in (Hansen, 1994; Bellinger et al., 2021; Nam et al., 2021; Krale et al., 2023); the last two references refer to this formulation as an Action-Contingent-Noiseless-Observable MDP, or ACNO-MDP. (Hansen, 1994) proposes a truncation-based approximation analogous to that in Section 4.2, except they provide to approximation guarantees. (Nam et al., 2021) and (Bellinger et al., 2021) focus on reinforcement learning (RL) (as opposed to the planning problem considered here). Specifically, Nam et al. focuses on developing RL algorithms for a fixed-horizon setting using the generic POMDP solver POMCP (Silver & Veness, 2010). On the other hand, Bellinger et al. adapts Q-learning for this setting by utilizing a statistical state estimator to achieve a “higher costed return” – for every non-state-sensing action, the subsequent state is simply sampled from the belief distribution. An $\epsilon$-greedy action is then taken based on the sampled state to update the Q-table, without leveraging any structure of the belief distribution while choosing the action. (Krale et al., 2023) proposes a policy improvement heuristic referred to as ATM and devises an RL algorithm to learn this heuristic; we contrast the heuristic proposed here to the ATM heuristic in Section 3, and also in our numerical case study in Section 5. Note that none of the above-mentioned papers focuses on the planning problem in a manner that exploits the specific structure of the MDP and from the standpoint of provable optimality/suboptimality guarantees.
A related formulation is considered in (Armstrong-Crews & Veloso, 2007), which treats “sensing” as a distinct action and applies a discount factor for its cost at each step a sensing action is taken. Aside from this distinction in the problem formulation, the JIV algorithm proposed in this paper is conceptually similar to the ATM heuristic proposed in (Krale et al., 2023).
Finally, another related formulation is analysed in (Reisinger & Tam, 2024); here, if the agent decides not to sense the state in any epoch, it is constrained to play the same action as in the previous epoch.
@@ 2 Problem Formulation
  • $\mathcal{S}=\{1,2,\ldots,|\mathcal{S}|\}$ denotes the (finite) state space,
  • $A=\{1,2,\ldots,|A|\}$ denotes the (finite) action space,
  • $\mathcal{T}$ denotes the transition function (i.e., $\mathcal{T}(s,a,s^{\prime})$ denotes the probability of transitioning to state $s^{\prime}$ on taking action $a$ in state $s$),
  • $\mathcal{C}$ denotes the cost function (i.e., $\mathcal{C}(s,a)$ is the cost associated with taking action $a$ in state $s$),
  • $\alpha\in(0,1)$ denotes the discount factor.
With some abuse of notation, for $a\in A,$ we use $\mathcal{T}(a)$ and $\mathcal{C}(a)$ to denote, respectively, the $|\mathcal{S}|\times|\mathcal{S}|$ transition probability matrix, and the $|\mathcal{S}|\times 1$ (column) vector of costs, associated with the action $a.$ [^1] Implicit in this notation is the assumption that any action $a\in A$ can be taken in any state $s\in\mathcal{S}.$ Denoting the state at time $t$ by $X_{t},$ and the action at time $t$ by $A_{t},$ there is a well-established theory for characterizing and computing the optimal policy that minimizes the expected discounted cost
@@
see (Puterman, 2014) and  (Ross, 1992). We use $V^{*}$ and $Q^{*}$ to denote, respectively, the optimal value function and the optimal action-value function, corresponding to $\mathcal{M}.$ As is convention, we treat $V^{*}$ to be an $|\mathcal{S}|\times 1$ column vector, and $Q^{*}$ to be a $|\mathcal{S}|\times|A|$ matrix.
MDP with state sensing cost: We now incorporate a positive state sensing cost $k$ to the above baseline MDP. Formally, the protocol for the interaction between the agent and the environment is as follows: At time $t\geq 0,$
  • Agent takes action $A_{t},$ and commits to either sensing the state at the next time step, or not; in the former case, it is said to have made a sensing action, and in the latter case, it is said to have made a blind action.
  • Agent incurs cost $\mathcal{C}(X_{t},A_{t}),$ and an additional sensing cost $k$ in case it made a sensing action.
  • The next state $X_{t+1}$ gets chosen randomly as per $\mathcal{T}(X_{t},A_{t},\cdot).$ In case the agent made a sensing step, then $X_{t+1}$ is revealed to it. If the agent made a blind step (thereby ‘saving’ on the sensing cost), then $X_{t+1}$ is not revealed to it.
We assume that the agent knows its initial state $X_{0}.$ Note that if the agent chooses a blind action at time $t,$ it must make its next action $A_{t+1}$ without precise knowledge of the state $X_{t+1}$ of the environment. The goal of the agent is to minimize its expected discounted cost (including the state sensing cost), i.e.,
@@
In the remainder of this section, we formulate the above sequential decision problem as an MDP $\mathcal{M}_{k}$ with a countably infinite state space. In the following, we refer to states in the baseline MDP (i.e., the elements of $\mathcal{S}$) as ‘root states.’ The state space of $\mathcal{M}_{k}$ is defined as
@@
Here, the state variable corresponds to the most recently sensed (root) state, along with the string of blind actions taken thereafter. Note that each state $\tilde{s}\in\mathcal{S}_{\infty}$ is associated with a belief distribution $\mathcal{B}(\tilde{s})\in\mathbb{R}^{1\times|\mathcal{S}|}$ over the set of root states. Specifically, for $\tilde{s}=(s,a_{1},a_{2},\ldots,a_{n}),$ where $s\in\mathcal{S}$ and $a_{i}\in A$ for $1\leq i\leq n,$
@@
where $e_{s}$ denotes the unit row vector with the $s^{\text{th}}$ entry being one. By convention, for $\tilde{s}=s\in\mathcal{S},$ (i.e., right after a sensing action), $\mathcal{B}(\tilde{s})=e_{s}.$
Next, the action space $\mathcal{A}$ for $\mathcal{M}_{k}$ is defined as
@@
where the second component of the action captures the decision of whether or not to sense the state at the next time step. Note that $\mathcal{A}$ is finite. We also write $\mathcal{A}=\mathcal{A}_{s}\cup\mathcal{A}_{b},$ where $\mathcal{A}_{s}=A\times\{sense\}$ denotes the set of sensing actions, and $\mathcal{A}_{b}=A\times\{blind\}$ the set of blind actions.
@@
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,blind),\tilde{s}_{2})$ =$=\begin{cases}1,&\text{for }\tilde{s}_{2}=(\tilde{s}_{1},a)\\ 0,&\text{otherwise}\end{cases}$
@@
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2})$ =$=\begin{cases}0,&\text{for }\tilde{s}_{2}\notin\mathcal{S}\\ \mathcal{B}(\tilde{s})\mathcal{T}(a)e_{\tilde{s}_{2}}^{T},&\text{for }\tilde{s}_{2}\in\mathcal{S}\end{cases}$
@@
The MDP $\mathcal{M}_{k},$ which captures opportunistic state sensing and a sensing cost, can now be defined formally using $(\mathcal{S}_{\infty},\mathcal{A},\mathcal{T}_{\infty},\mathcal{C}_{\infty},\alpha)$. Given that the state space of this MDP is countable and its action space is finite, there exists an optimal stationary policy (see (Puterman, 2014) and  (Ross, 1992)); however, this optimal policy is not amenable to an exact computation given the infinite state space. In Section 3, we propose a heuristic based on selective policy improvement, and in Section 4, we propose iterative schemes for computing an optimal (or near-optimal) policy via a truncation of the state space.
3 Heuristic Algorithm
In this section, we introduce the Selective Policy Improvement (SPI) heuristic for our opportunistic state sensing MDP. This heuristic is stated formally as Algorithm 1.
@@ 3 Heuristic Algorithm
Setting up our notation, define $V_{\mathcal{M}_{k}}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times 1}$ as the value function column vector corresponding to the policy $\pi$ for the root states. The notation $\pi^{\prime}\overset{\mathcal{S}}{\succeq}\pi$ indicates that the vector $V_{\mathcal{M}_{k}}^{\pi^{\prime}}$ is element-wise less than or equal to $V_{\mathcal{M}_{k}}^{\pi}$. Finally, $\mathsf{\bf max}\ x$ (respectively, $\mathsf{\bf min}\ x$) for a vector $x$ denotes its maximum (respectively, minimum) entry.
@@
The SPI heuristic, initialized with a certain initial policy $\pi_{init}$ and hyperparameters $\delta$ and maxsteps, calls the $\mathsf{PolicyUpdate}$ routine (see Algorithm 2) repeatedly, until the value function improvement falls below a prescribed threshold $\delta.$ The $\mathsf{PolicyUpdate}$ routine in turn improves upon its input policy $\pi_{ref}$ as follows. For each root state, it tries to find a blind action sequence having length at most maxsteps, which improves upon $\pi_{ref}.$
@@
While the set of all such blind action paths can be quite large, $\mathsf{PolicyUpdate}$ searches this space selectively for computational tractability. Specifically, for a root state $s,$ it considers the blind action trajectory $(a_{1},a_{2},\ldots,a_{n})$ only when for each $i,$ $(a_{1},a_{2},\ldots,a_{i})$ is an improvement over $(a_{1},a_{2},\ldots,a_{i-1})$ followed by the optimal sensing action. This allows for an efficient (and greedy) search for an improving blind action trajectory.
@@
Note that $V_{MS}(\mathcal{B}(\tilde{s}),\bar{V})$ denotes the value associated with playing the optimal sensing action under belief $\mathcal{B}(\tilde{s})$ with terminal values $\bar{V}$ (MS here stands for myopic sensing). $\pi_{MS}$ denotes the corresponding optimal sensing action. Importantly, note that the SPI heuristic only performs policy evaluations over the root states of $\mathcal{M}_{k}.$
@@
1:  $\pi^{\prime}\leftarrow\pi_{init}$
2:  $\pi_{\text{improv}}\leftarrow\textsc{PolicyUpdate}(\pi^{\prime},\text{\emph{maxsteps}})$
@@
Finally, we note that in our numerical case studies, we initialize the SPI heuristic with the Always Sense (AS) policy, which takes the optimal sensing action in each belief state. The value function of the AS policy is, for $\tilde{s}\in\mathcal{S}_{\infty},$
$V_{AS}(B(\tilde{s}))$ $=\min_{a\in A}\bigl{(}B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)V^{*}\bigr{)}+\frac{k}{1-\alpha}$
@@
Here, $\mathsf{\bf argmin}\ x$ for a row vector $x$ denotes the column index corresponding to its minimum entry. It is important to note that AS policy $\pi_{AS}$ agrees with the optimal policy $\pi^{*}$ associated with the baseline MDP $\mathcal{M}$ over root states. In Section 4.1, we show that the AS policy is also optimal for $\mathcal{M}_{k}$ when the sensing cost $k$ is small.
It is instructive at this point to compare the SPI heuristic with the ATM heuristic proposed in (Krale et al., 2023). The latter is more restricted in its search for good blind actions; in any belief state $\tilde{s},$ it seeks to improve upon $\pi_{AS}$ by comparing the actions $(\pi_{AS}(\tilde{s}),\textit{blind})$ and $(\pi_{AS}(\tilde{s}),\textit{sense}).$

### v3

@@ Abstract
In many practical sequential decision-making problems, tracking the state of the environment incurs a sensing/computation cost. In these settings, the agent’s interaction with its environment includes the additional component of deciding when to sense the state, in a manner that balances the value associated with optimal (state-specific) actions and the cost of sensing. We formulate this as an expected discounted cost Markov Decision Process (MDP), wherein the agent incurs an additional cost for sensing its next state, but has the option to take actions while remaining ‘blind’ to the system state. We pose this problem as a classical discounted cost MDP with an expanded (countably infinite) state space. While computing the optimal policy for this MDP is intractable in general, we derive lower bounds on the optimal value function, which allow us to bound the suboptimality gap of any policy. We also propose a computationally efficient algorithm SPI, based on policy improvement, which in practice performs close to the optimal policy. Finally, we benchmark against the state-of-the-art via a numerical case study.
@@ 1 Introduction
Markov Decision Processes (MDPs) provide a powerful framework for modeling sequential interactions between an agent and an adaptive environment that ‘responds’ to the agent’s actions. In the classical MDP formulation, at each time step $t$, the agent sees the environment’s state, denoted $X_{t}$, and selects an action $A_{t}$. This action generates a feedback signal (cost) determined by the state-action pair $(X_{t},A_{t})$, and triggers a random, action-dependent Markovian transition in the state. However, in many applications, ‘seeing’ the current state involves a cost. For example:
$\bullet$ Healthcare: Monitoring a patient’s state during an ongoing intervention, such as white blood cell counts in anti-HIV therapy or lab tests for ICU patients, incurs monetary or delay costs (Ernst et al., 2006).
$\bullet$ Mobile Applications: Sensing a user’s location, motion, or environment on mobile devices incurs energy costs, which must be balanced with user experience goals (Wang et al., 2010).
$\bullet$ Wireless Sensor Networks: Sensing the network state might require turning on battery-operated (and therefore energy-constrained) sensors.
$\bullet$ Distributed Sensing: Aggregating sensor data to determine system state involves computational costs. $\bullet$ Remote Surveillance: Transmitting state information (in the form of images/video) to a controller incurs communication costs.
$\bullet$ Robotics: In applications where an autonomous robot is engaged in a task, sensing the surroundings might induce a latency (cost) in task completion.
This motivates integrating the cost of state sensing and a model for opportunistic state sensing into the MDP framework. However, such integration introduces technical challenges. In particular, if the agent chooses not to sense the system state (while continuing to interact with the environment), it is left with a belief distribution over the system state. Incorporating this belief into the MDP formulation induces a state space explosion, which makes the decision problem intractable.
@@
While the ‘augmented’ MDP admits a stationary Markov policy, it is computationally intractable given the infinite state space. The goal of this paper is to design tractable, near-optimal algorithms for solving it. Our key contributions are as follows:
$\bullet$ We propose SPI, a policy-improvement-based algorithm that selectively searches for improving blind action sequences and is near-optimal in practice.
@@
$\bullet$ We utilize the truncated MDPs to derive lower bounds on the optimal value function, enabling computation of the suboptimality gap for any policy.
@@
$\bullet$ We derive a computable state-sensing cost threshold, below which always sensing is optimal.
$\bullet$ Finally, we conduct an extensive numerical case study, including a real-world healthcare application, demonstrating that SPI consistently outperforms state-of-the-art POMDP solvers under diverse sensing costs, while maintaining reasonable compute time.
@@
It is worth noting that our formulation can be posed as a Partially Observable Markov Decision Processes (POMDP) and indeed we benchmark our methods against state-of-the-art POMDP solvers (see Section 5). However, in doing so, one loses the specific problem structure that arises in the opportunistic state sensing formulation, which we seek to exploit here for computational tractability. Since POMDPs are known to be intractable in general (Papadimitriou and Tsitsiklis, 1987), leveraging this structure is crucial. Specifically, our algorithms and analytical results exploit the specific structure of this formulation, namely that the uncertainty in the belief distribution can be collapsed entirely by paying a sensing cost, allowing a good policy to go ‘blind’ until the cost induced by state uncertainty is outweighed by the sensing cost. A generic MDP/POMDP solver would not be able to exploit this structure. Moreover, our algorithmic design and proof arguments rely on classical MDP tools–policy improvement and the Bellman operator–whose POMDP analogues are significantly weaker.
@@
Related Literature. Formulations equivalent to ours have been analysed in (Hansen, 1994; Bellinger et al., 2021; Nam et al., 2021; Krale et al., 2023); the last two references refer to this formulation as an Action-Contingent-Noiseless-Observable MDP, or ACNO-MDP. Hansen (1994) proposes a truncation-based approximation analogous to that in Section 4.2, except they provide no approximation guarantees. (By contrast, Chen and Liew (2025) develops a truncation-based approach with theoretical guarantees for a simpler, related setting where the state uncertainty is exogenous.) Nam et al. (2021); Bellinger et al. (2021) focus on reinforcement learning (RL) (as opposed to the planning problem considered here). Specifically, Nam et al. (2021) focuses on developing RL algorithms for a fixed-horizon setting using the generic POMDP solver POMCP (Silver and Veness, 2010). On the other hand, Bellinger et al. (2021) adapts Q-learning for this setting by utilizing a statistical state estimator to achieve a “higher costed return” – for every non-state-sensing action, the subsequent state is simply sampled from the belief distribution. An $\epsilon$-greedy action is then taken based on the sampled state to update the Q-table, without leveraging any structure of the belief distribution while choosing the action. Krale et al. (2023) proposes a policy improvement heuristic referred to as ATM and devises an RL algorithm to learn this heuristic; we contrast the algorithm proposed here to the ATM heuristic in Section 3, and also in our numerical case study in Section 5. Note that none of the above-mentioned papers focuses on the planning problem in a manner that exploits the specific structure of the MDP and from the standpoint of provable optimality/suboptimality guarantees. A related formulation is considered in Armstrong-Crews and Veloso (2007), which treats “sensing” as a distinct action and applies a discount factor for its cost at each step a sensing action is taken. Aside from this distinction in the problem formulation, the JIV algorithm proposed in this paper is conceptually similar to the ATM heuristic proposed in Krale et al. (2023). A similar framework also appears in Treven et al. (2024), but for continuous-time problems with emphasis on learning rather than planning. Finally, another related formulation is analysed in Reisinger and Tam (2024); here, if the agent decides not to sense the state in any epoch, it is constrained to play the same action as in the previous epoch.
@@ 2 Problem Formulation
$\bullet$ $\mathcal{S}=\{1,2,\ldots,|\mathcal{S}|\}$ denotes the (finite) state space,
$\bullet$ $A=\{1,2,\ldots,|A|\}$ denotes the (finite) action space,
$\bullet$ $\mathcal{T}$ denotes the transition function (i.e., $\mathcal{T}(s,a,s^{\prime})$ denotes the probability of transitioning to state $s^{\prime}$ on taking action $a$ in state $s$),
$\bullet$ $\mathcal{C}$ denotes the cost function (i.e., $\mathcal{C}(s,a)$ is the cost associated with taking action $a$ in state $s$),
$\bullet$ $\alpha\in(0,1)$ denotes the discount factor.
With some abuse of notation, for $a\in A,$ we use $\mathcal{T}(a)$ and $\mathcal{C}(a)$ to denote, respectively, the $|\mathcal{S}|\times|\mathcal{S}|$ transition probability matrix, and the $|\mathcal{S}|\times 1$ (column) vector of costs, associated with the action a.a. [^1] Implicit in this notation is the assumption that any action $a\in A$ can be taken in any state $s\in\mathcal{S}.$ Denoting the state at time $t$ by $X_{t},$ and the action at time $t$ by $A_{t},$ there is a well-established theory for characterizing and computing the optimal policy that minimizes the expected discounted cost
@@
see Puterman (2014) and  Ross (1992). We use $V^{*}$ and $Q^{*}$ to denote, respectively, the optimal value function and the optimal action-value function, corresponding to $\mathcal{M}.$ As is convention, we treat $V^{*}$ to be an $|\mathcal{S}|\times 1$ column vector, and $Q^{*}$ to be a $|\mathcal{S}|\times|A|$ matrix.
MDP with state sensing cost: We now incorporate a state sensing cost $k\hskip-2.0pt>\hskip-2.0pt0$ to the above baseline MDP. The interaction protocol between the agent and environment at each time step $t\hskip-2.0pt\geq\hskip-2.0pt0$ is as follows:
$\bullet$ The agent takes action $A_{t}$ and commits either to sensing the state at the next step (a sensing action) or not (a blind action).
$\bullet$ Agent incurs cost $\mathcal{C}(X_{t},A_{t}),$ and an additional sensing cost $k$ in case it made a sensing action.
$\bullet$ The next state $X_{t+1}$ is drawn according to $\mathcal{T}(X_{t},A_{t},\cdot)$. If the agent made a sensing action, then $X_{t+1}$ is revealed; if it made a blind action, then $X_{t+1}$ remains hidden (saving on the sensing cost $k$).
We assume the agent knows its initial state $X_{0}$. If it chooses a blind action at time $t$, then its next action $A_{t+1}$ must be taken without precise knowledge of $X_{t+1}$. The agent’s objective is to minimize expected discounted cost (including the state sensing cost), i.e.,
@@
We formulate the sequential decision problem as an MDP $\mathcal{M}_{k}$ with a countably infinite state space. For clarity, we refer to states in the baseline MDP (elements of $\mathcal{S}$) as ‘root states.’ The state space of $\mathcal{M}_{k}$ is
@@
Here, the state variable corresponds to the most recently sensed (root) state, along with the string of blind actions taken thereafter. Each state $\tilde{s}\in\mathcal{S}_{\infty}$ is associated with a belief distribution $\mathcal{B}(\tilde{s})\in\mathbb{R}^{1\times|\mathcal{S}|}$ over the set of root states. Specifically, for $\tilde{s}=(s,a_{1},\ldots,a_{n}),$ where $s\in\mathcal{S}$ and $a_{i}\in A$ for $1\leq i\leq n,$
@@
where $e_{s}$ denotes the unit row vector with the $s^{\text{th}}$ entry being one. By convention, for $\tilde{s}=s\in\mathcal{S},$ (i.e., right after a sensing action), $\mathcal{B}(\tilde{s})=e_{s}.$ Next, the action space $\mathcal{A}$ for $\mathcal{M}_{k}$ is defined as
@@
where the second component of the action captures the decision of whether or not to sense the state at the next time step. Thus $\mathcal{A}$ is finite; we also write $\mathcal{A}=\mathcal{A}_{s}\cup\mathcal{A}_{b},$ where $\mathcal{A}_{s}=A\times\{sense\}$ are sensing actions, and $\mathcal{A}_{b}=A\times\{blind\}$ blind actions.
@@
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,blind),\tilde{s}_{2})$ $=\begin{cases}1,&\text{for }\tilde{s}_{2}=(\tilde{s}_{1},a)\\ 0,&\text{otherwise}\end{cases}$
@@
$\mathcal{T}_{\infty}(\tilde{s}_{1},(a,sense),\tilde{s}_{2})$ $=\begin{cases}0,&\text{for }\tilde{s}_{2}\notin\mathcal{S}\\ \mathcal{B}(\tilde{s})\mathcal{T}(a)e_{\tilde{s}_{2}}^{T},&\text{for }\tilde{s}_{2}\in\mathcal{S}\end{cases}$
@@
The MDP $\mathcal{M}_{k},$ modeling opportunistic state sensing with a sensing cost, is defined using $(\mathcal{S}_{\infty},\mathcal{A},\mathcal{T}_{\infty},\mathcal{C}_{\infty},\alpha)$. Since the state space is countable and the action space finite, there exists an optimal stationary policy  (Puterman, 2014; Ross, 1992); however, the exact computation of the optimal policy is infeasible due to the infinite state space. In Section 3, we propose an algorithm based on selective policy improvement, and in Section 4, iterative schemes for computing an optimal (or near-optimal) policy via state-space truncation together with a lower bound on the optimal value function, allowing us to quantify the suboptimality gap of any policy.
It is worth noting that, while POMDPs can be reformulated as MDPs over belief states (i.e., belief MDPs), such a transformation leads to an uncountably infinite state space, as the belief space forms a continuous (n1)(n-1)-simplex. In contrast, our formulation—an MDP with sensing costs—can be viewed as a special case of a POMDP that admits a countable state-space representation. This structural property makes our problem significantly more tractable than general POMDPs.
3 Selective Policy Improvement
@@
In this section, we introduce the Selective Policy Improvement (SPI) algorithm for the opportunistic state sensing MDP. The algorithm is stated formally as Algorithm 1. For our notation, let $V_{\mathcal{M}_{k}}^{\pi}\in\mathbb{R}^{|\mathcal{S}|\times 1}$ denote the value function column vector of policy $\pi$ for the root states. The notation $\pi^{\prime}\overset{\mathcal{S}}{\succeq}\pi$ indicates that the vector $V_{\mathcal{M}_{k}}^{\pi^{\prime}}$ is element-wise less than or equal to $V_{\mathcal{M}_{k}}^{\pi}$. Finally, $\mathsf{\bf max}\ x$ (respectively, $\mathsf{\bf min}\ x$) for a vector $x$ denotes its maximum (respectively, minimum) entry.
@@
SPI is initialized with a policy $\pi_{init}$ and hyperparameters $\delta$ and maxsteps. It iteratively applies the PolicyUpdate routine (Algorithm 2) until the improvement in the value function falls below $\delta$. The PolicyUpdate routine improves a reference policy $\pi_{ref}$ as follows:
@@
For each root state $s,$ it searches for an ‘improving’ sequence of blind actions of length at most maxsteps, which lowers the action-value function at $s,$ relative to the reference policy $\pi_{\text{ref}}$. While the set of all such blind action sequences can be quite large, $\mathsf{PolicyUpdate}$ searches this space selectively for computational tractability. Specifically, for a root state $s,$ $\pi^{\prime}$ considers the candidate blind action sequence $(a_{1},\ldots,a_{n})$ only when for each $i,$ $(a_{1},\ldots,a_{i})$ is an improvement over $(a_{1},\ldots,a_{i-1}),$ followed by the optimal sensing action. This allows for an efficient (and greedy) search for an improving blind action trajectory.
@@
[…]
@@
[…]
@@
Finally, in our numerical case studies, we initialize the SPI algorithm with the Always Sense (AS) policy, which selects the optimal sensing action at each belief state. The corresponding value function of the AS policy, for $\tilde{s}\in\mathcal{S}_{\infty},$ is given by
$)\displaystyle V_{AS}(B(\tilde{s}$)) $=\min_{a\in A}\bigl(B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)V^{*}\bigr)+\frac{k}{1-\alpha}$
@@
Here, $\mathsf{\bf argmin}\ x$ for a row vector $x$ denotes the column index of its minimum entry. It is important to note that AS policy $\pi_{AS}$ agrees with the optimal policy $\pi^{*}$ associated with the baseline MDP $\mathcal{M}$ over root states. In Section 4.1, we show that the AS policy is also optimal for $\mathcal{M}_{k}$ when the sensing cost $k$ is small.
3.1 Step-by-Step Walkthrough of $\mathsf{Policy\ Update}$
Starting at every root state $s$, $\mathsf{PolicyUpdate}$ initializes the candidate policy $\pi^{\prime}$ with $\pi_{ref}$. It then searches for a sequence of actions to update $\pi^{\prime}$ and improve upon $\pi_{ref}$ at the specific root state $s$. For the given root state, the output policy $\pi_{o}$ plays the sequence of action that performs better at that root state.
Lines 9–10. For a given belief state $\tilde{s}$ (initialized by the root state $s$), the algorithm identifies the blind action $a_{blind}$ as follows: For each blind action $(a,blind)\in\mathcal{A}_{b}$, it computes the cumulative cost, which consists of:
$\bullet$ Executing the blind action $(a,blind)$, incurring an immediate cost $\mathcal{B}(\tilde{s})\mathcal{C}(a)$.
$\bullet$ Transitioning to the belief state $(\tilde{s},a_{blind})$ and obtaining the subsequent return $V_{MS}(\mathcal{B}(\tilde{s})\mathcal{T}(a),V^{\pi_{ref}}_{\mathcal{M}_{k}}).$
Note that this subsequent cumulative cost of $V_{MS}(\mathcal{B}(\tilde{s})\mathcal{T}(a),V^{\pi_{ref}}_{\mathcal{M}_{k}})$is obtained by first executing the sensing action given by $\pi_{MS}(\mathcal{B}(\tilde{s})\mathcal{T}(a),V^{\pi_{ref}}_{\mathcal{M}_{k}})$ at the belief state $(\tilde{s},a_{blind})$, and then following the reference policy $\pi_{ref}$ thereafter. The minimum cumulative cost over all blind actions is denoted $V_{blind}$, with the corresponding minimizing action denoted $a_{blind}$.
Lines 11–17. If the cumulative cost of playing the optimal sensing action at $\tilde{s}$ given by $V_{MS}(\mathcal{B}(\tilde{s}),V^{\pi_{ref}}_{\mathcal{M}_{k}}),$ is lower than $V_{blind}$, the algorithm terminates the search and updates the candidate policy $\pi^{\prime}$ with the obtained sequence of actions, including the final sensing action. Otherwise (lines 15-17), the process repeats (up to a maximum of maxstepsmaxsteps) for the subsequent belief state $(\tilde{s},a_{blind})$ obtained by executing the action $(a_{blind},blind)$ on state $s$.
Lines 19–23. The value function $V_{\mathcal{M}_{k}}^{\pi^{\prime}}(s)$ of the policy $\pi^{\prime}$ at $s$, which evaluates the cumulative cost of the candidate action sequence, is compared to that of the reference policy $\pi_{ref}$. The output policy $\pi_{o}$ then adopts the sequence of actions starting from $s$ that corresponds to the policy with the lower value function.
This process is repeated for all root states, constructing the final output policy $\pi_{o}$ by aggregating the improved action sequences for each root state where the value of $\pi^{\prime}$ is lower than that of the reference policy $\pi_{ref}$.
It is instructive to compare SPI with the ATM heuristic proposed in Krale et al. (2023). The latter is more restricted in its search for good blind actions; in any belief state $\tilde{s}$, it seeks to improve upon $\pi_{AS}$ by comparing the actions $(\pi_{AS}(\tilde{s}),\textit{blind})$ and $(\pi_{AS}(\tilde{s}),\textit{sense}).$ For a detailed analysis of the ATM heuristic, we refer the reader to Section D.